i dont know how to find the perimeter of a rectangle with the long side 2w + 3 and the short side w. the rectangle has 4 90 degree angles.
thanks!!

Found 2 solutions by rothauserc, solver91311:
Answer by rothauserc(4718)

The following is the stated problem
i dont know how to find the perimeter of a rectangle with the long side 2w + 3 and the short side w. the rectangle has 4 90 degree angles.
The perimeter (P) of a rectangle is twice the length(l) plus twice the width(w).
We have l = 2w+3, therefore
P = 2(2w+3) + 2w
P = 4w+6+2w = 6w+6 = 6(w+1)

Answer by solver91311(24713)



If you are expecting to assign a specific number to the value of the perimeter of your rectangle, you are doomed to disappointment. The best you can do is create an expression in that represents your perimeter. The perimeter of a rectangle is two times the length plus two times the width, thus:



But since you know that , you can substitute and so define the perimeter as a function of the width:



You will need to simplify the expression yourself. And, just as an 'oh by the way', it darned well better have four right angles, otherwise you had better stop calling it a rectangle.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism